# Far from Perfect: Quantum Error Correction with (Hyperinvariant) Evenbly Codes

### v1

@@ Abstract
We introduce a new class of qubit codes that we call Evenbly codes, building on a previous proposal of hyperinvariant tensor networks. Its tensor network description consists of local, non-perfect tensors describing CSS codes interspersed with Hadamard gates, placed on a hyperbolic $\{p,q\}$ geometry with even $q\geq 4$, yielding an infinitely large class of subsystem codes. We construct an example for a $\{5,4\}$ manifold and describe strategies of logical gauge fixing that lead to different rates $k/n$ and distances $d$, which we calculate analytically, finding distances which range from $d=2$ to $d\sim n^{2/3}$ in the ungauged case. Investigating threshold performance under erasure, depolarizing, and pure Pauli noise channels, we find that the code exhibits a depolarizing noise threshold of about $19.1\%$ in the code-capacity model and $50\%$ for pure Pauli and erasure channels under suitable gauges. We also test a constant-rate version with $k/n=0.125$, finding excellent error resilience (about $40\%$) under the erasure channel. Recovery rates for these and other settings are studied both under an optimal decoder as well as a more efficient but non-optimal greedy decoder. We also consider generalizations beyond the CSS tensor construction, compute error rates and thresholds for other hyperbolic geometries, and discuss the relationship to holographic bulk/boundary dualities. Our work indicates that Evenbly codes may show promise for practical quantum computing applications.
I Introduction
@@
In this work, we explore a recent proposal for hyperinvariant tensor network (HTN) codes [24] and show that it encompasses simple qubit-level codes; we dub these new codes Evenbly codes after the author of the original HTN paper [25]. An essential property of these tensor network codes is that they do not require perfect tensors, and can thus be built from gates that mediate less-than-maximal entanglement. Our specific construction incorporates seed tensors from $\llbracket q,1,2\rrbracket_{2}$ CSS codes placed on the vertices of a hyperbolic tiling, along with Hadamard gates on its edges; this can be generalized to any even-$q$ hyperbolic tiling with Schläfli symbols $\{p,q\}$, giving rise to a new, infinite family of holographic subsystem codes. We methodically show that the resulting code structure upholds the isometric constraints typified by the HTN ansatz [26, 25], and additionally show how to partition the logical subspace as a subsystem code. The distance and rate of the code is analyzed, and variants with a rate up to $R=\sqrt{3}/2$ and a distance scaling up to $d(n)\sim n^{2/3}$ can be constructed. Moreover, we investigate the erasure threshold of the zero-rate $\{5,4\}$ Evenbly code using two types of decoders: a quadratically-scaling greedy reconstruction algorithm and an optimal recovery algorithm based on Gaussian elimination, which exhibits a cubic time complexity [9, 27]. We find that the greedy reconstruction algorithm faithfully provides a lower bound for the threshold (approximately $40\%$), whereas the Gaussian-elimination algorithm obtains a threshold of $49.95\pm 0.01\%$ in the $Z$ gauge. A constant-rate $k/n=0.125$ variant of the Evenbly code is tested as well, and it is shown to achieve a threshold of about $44\%$ in the $Y$ gauge for the central logical qubit, and more than $39\%$ in the $Z$ gauge. In both the zero- and constant-rate case, the $X$ gauge corresponds to a product state embedding of the ungauged logical qubits on the boundary, preserving the transversal weight-$2$ gates of the $\llbracket 4,1,2\rrbracket_{2}$ seed code. We also study the effects of depolarizing and pure Pauli noise on the zero-rate $\{5,4\}$ Evenbly code using the integer-optimization decoder from [28], proving the existence of a threshold at $19.1\pm 0.94\%$ for the former, and around $50\%$ for the latter, ostensibly surpassing the well-known zero-rate hashing bound from random coding theory [29, 30, 31]. Other constructions are possible, and we provide several examples that utilize seed tensors which are neither CSS nor planar 2-uniform, as well as evaluating the erasure thresholds under greedy decoding for several even-$q$ generalizations of the Evenbly code discussed in the main text.
@@
This paper is organized as follows. We start with a synopsis on tensor networks and holographic quantum codes (Section II.1) and subsequently introduce hyperinvariant tensor networks (HTN) as an ansatz (Section II.2). Next, we construct Evenbly codes (Section III), commencing first with the introduction of 1-uniform (i.e. GHZ) states to the vertex tensors of the original HTN ansatz (Section III.1), and following subsequently with the construction of the qubit-level Evenbly code using planar 2-uniform states (Section III.2). We further show that our construction admits a subsystem structure, in line with previous holographic quantum codes, and we describe how to partition the logical space and perform gauge fixing for the code (Section III.3). Proceeding, we investigate the rate and distance scaling for the code, and as a result construct a finite-rate version of the Evenbly code (Section III.4); the threshold properties are investigated for a zero-rate Evenbly code (Section III.5) and a constant-rate variant in Section III.6; in the process, we explain how the $X$ gauge choice in the Evenbly code allows for weight-2 transversal logical operations to emerge, and we show that very high thresholds emerge in several of the other gauge choices. Finally we report depolarizing and pure Pauli noise threshold results (Section III.7), which are found to be competitive with known code-capacity thresholds for topological codes [6]. The implications of our findings are discussed and concluding comments are provided in Section IV. In the appendices, we provide additional information on how to build multipartite maximally-entangled (MME) states (Appendix A); an example construction of an Evenbly code using an AME graph state (Appendix B); Several additional constructions on generalized $\{p,q\}$ tilings and and an accompanying greedy threshold analysis (Appendix C); several seed-tensor constructions that are neither CSS nor planar 2-uniform states, yet still can be used to create new Evenbly codes (Appendix D); and lastly, details on the three different decoder strategies employed in this paper (Appendix E).
II Background
II.1 Tensor Networks & Holographic Codes
@@
Holographic codes typically exhibit uberholography, i.e., only operators on a fractal subset of the boundary are needed in order to reconstruct logical operators in the bulk [40]. Uberholography is a special case of operator-algebra quantum-error correction [41], in which a subalgebra representing all possible logical operators in a region is defined using non-local physical boundary operators. HaPPY-style holographic codes have been shown to exhibit high thresholds under both erasure and depolarizing noise models [28, 14, 13]. In the spirit of investigating holographic dualities, it has been shown that HaPPY-style holographic codes do not reproduce many of holography’s known features at finite $N$ (i.e., under gravitational corrections) [42, 43]. Such features include non-trivial entanglement spectra of reduced density matrices, state-dependent bulk-reconstruction, and corrections to the entanglement entropy, whose recovery requires tensor network codes with non-perfect tensors [11, 12, 24]. The quantum code construction that we propose in this paper, in addition to its novel error correction properties, are thus relevant from the perspective of holography as well; we address these attributes in our companion paper [24].
II.2 Hyperinvariant Tensor Networks
Figure 1: Codes from a hyperinvariant tensor network (HTN). (a) Evenbly’s HTN setup from [25], composed of $A$ and $B$ tensors on a hyperbolic $\{p,q\}$ tiling (here for $p=5,q=4$) that respect the symmetries of the infinitely extended tiling. Open black edges represent a state on the physical boundary qubits, while connected black edges are tensor network contractions. (b) The extension of the HTN into a quantum error correction code proposed in [24] (dubbed Evenbly code), in which the $q$-leg $A$ tensors are replaced by $q{+}1$-leg $A^{\prime}$ tensors, with the additional leg (red dot) representing a logical “bulk” qubit. The tensor network forms the encoding isometry of this bulk-to-boundary code.
@@ II.2 Hyperinvariant Tensor Networks
Figure 2: Isometry constraints for the HTN defined on a $\{p,q\}$ tiling. We color-code the relevant tensors as follows: light-/dark-gray tensors represent $A$ and $A^{\dagger}$; light-/dark-blue tensors represent $B$ and $B^{\dagger}$; light-/dark-purple tensors depict the $3:1$ isometries $W$ and $W^{\dagger}$; and the $4:2$ isometries $U$ and $U^{\dagger}$ are shown in creme and light orange, respectively.
III Results
III.1 HTN Ansätze from 1-Uniform States
We begin by constructing first a simpler case of the original HTN ansatz of [25]; namely, we prepare a single boundary state rather than a bulk-to-boundary code, which permits us to analyze the tensor network structure more easily. We then show in Section III.2 that this simplified case easily generalizes to a new subclass of holographic quantum codes.
@@ III.1 HTN Ansätze from 1-Uniform States
    $A$ and $B$ fulfill isometry constraints involving tensors on single and neighboring vertices. The constraints are shown explicitly in Figure 2 for the $q>3$ case (the special $q=3$ case is covered in [25]).
@@
    $A$ is symmetric under cyclic index permutations, i.e., $A_{k_{1},k_{2},\dots,k_{q}}=A_{k_{q},k_{1},\dots,k_{q-1}}$.
@@
    $A$ is chosen as a non-perfect tensor, i.e., there exist bipartitions of its $q$ indices between which it does not mediate maximal entanglement.
@@
|$\ket{\text{GHZ}}=\frac{1}{\sqrt{2}}\left[\ket{0}^{\otimes q}+\ket{1}^{\otimes q}\right].$ (1)
@@
$\mathcal{G}_{1}$ $=X^{\otimes q},$ (2)
$\mathcal{G}_{k}$ $=I^{\otimes k-2}(ZZ)I^{\otimes q-k}\ ,$ (3)
where $k\in\{2,3,\dots,q\}$. As usual, we end up with $(n-k)=q$ generators for the full stabilizer group.
@@
$H=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix}\ .$ (4)
The isometry constraints from the remaining first criterion can be proven by using operator pushing: By applying the stabilizers from Equation 2, we can replace a Pauli operator acting on one of the tensor legs by an equivalent operator acting on other legs. Moving Pauli operators past a $B$ tensor involves exchanging $X\leftrightarrow Z$, as $XH=HZ$ and $ZH=HX$.
@@
Figure 3: Operator recovery on subsystems of the HTN with $A$ tensors (light gray) chosen to represent GHZ states and $B$ tensors (light blue) chosen as Hadamard gates. A) For the isometry $W$, applying the generators from Equation 2 removes a Pauli operator on the top leg while adding new ones on the bottom ones. As this is possible for the whole Pauli algebra (generated by X and Z), the tensors form an isometry from top to bottom legs, as expected from the 1-uniformity requirement. Note that pushing through a $B$ tensor swaps $X\leftrightarrow Z$. B) For the $U$ isometry, a Pauli operator can again be pushed to the bottom legs by applying the relation from Equation 2. Due to reflection symmetry, it is sufficient to consider operators acting on only one of the two top legs. These diagrams offer proof that the generalized isometry constraints shown in Figure 2 are upheld, thus allowing for the construction of a consummate HTN.
III.2 HTN qubit codes
@@
Now we generalize the results gained from Section III.1 and define Evenbly codes; i.e., HTNs with bulk logical degrees of freedom. Evenbly codes were first defined in [24] as tensor network codes that associate a logical qubit with each vertex in an HTN, a process of bulk implantation which replaces each $q$-leg $A$ tensor by a $(q+1)$-leg tensor $A^{\prime}$. In analogy with Definition 1, an Evenbly code is then defined as follows:
@@
Note that $A^{\prime}$ merely describes an error-detecting code with code distance $d=2$; this condition implicitly excludes perfect tensors, which would lead to codes with $d>2$ for $q>3$. Generally, finding solutions to the constraints in Figure 4 is much harder than for those in Figure 2 since the isometries $W^{\prime}$ and $U^{\prime}$ act on more input sites than $W$ and $U$.
[24] introduced only a single ququart example fulfilling these constraints; we will now show that our GHZ HTN construction can be extended to define a class of qubit codes whose $A^{\prime}$ tensors define Calderbank-Steane-Shor (CSS) codes. While we focus on this construction in the remainder of the paper, other qubit-level Evenbly codes can be constructed: In Appendices B and D, we introduce Evenbly codes whose seed tensors are not built from CSS codes, with $A^{\prime}$ tensors describing strictly 2-uniform or AME$(5,2)$ states. In contrast, the Evenbly code can be defined for any even $q\geq 4$, and its stabilizer generators defining the $A^{\prime}$ tensors of rank $(q+1)$ are given by
@@
|$\ket{\bar{0}}$ $=\frac{1}{\sqrt{2}}\left[\ket{0}^{\otimes q}+\ket{1}^{\otimes q}\right]\ ,$ (6a)
|$\ket{\bar{1}}$ $=\frac{1}{\sqrt{2}}\left[(\ket{0}\ket{1})^{\otimes\frac{q}{2}}+(\ket{1}\ket{0})^{\otimes\frac{q}{2}}\right]\ .$ (6b)
Both codeword states are invariant under cyclic permutation of the physical qubits, and hence the encoding isometry represented by $A^{\prime}$ is invariant under cyclic permutation of the physical indices. This code is closely related to the GHZ state construction with generators in Equation 2, which can be obtained by adding $\bar{Z}$ (defined below) as an additional generator to Equation 5. In fact, |$\ket{\bar{0}}$, |$\ket{\bar{1}}$, and any superposition of these two state vectors comprise a GHZ state in different bases, with each single site maximally entangled with the rest of the system.
The tensor $A^{\prime}$ is therefore partially 2-uniform when considering the logical and any one physical index, fulfilling the first of the Evenbly code constraints. In fact, it fulfills an even stronger condition: For any planar embedding of the logical index, the tensor describes a $q{+}1$-qubit state that is planar 2-uniform, i.e., any two qubits that are neighboring under this embedding are maximally entangled with the remainder [50]. One choice for the logical operators is given by
@@
Figure 5: Operator recovery on subsystems of the Evenbly code (Equation 5) via the isometry constraints. A)-B) For a single $A^{\prime}$ tensor, the logical $\bar{X}$ and $\bar{Z}$ operators (red) can be represented on $\frac{q}{2}$ and $2$ sites, respectively, following Equation 7. Note that X and Z operators are swapped when pushed through a $B$ tensor. For two adjacent $A^{\prime}$ tensors and an intermediate $B$ tensor, the $\bar{Z}$ operator on a given logical qubit can be represented on the physical qubits on one side. The $\bar{X}$ operator requires a representation on both sides, with an application of the stabilizers (Equation 5) removing any Pauli operators acting on contracted legs. By reflection symmetry, the same rules apply for logical operators acting on the second site. C)-D) Any operator acting on a single physical leg can also be pushed to the remaining ones by applying Equation 5. Together with A)-B), these rules are sufficient to prove the isometry conditions.
III.3 Evenbly Codes as Subsystem Codes

### v2

@@ Abstract
We introduce a new class of qubit codes that we call Evenbly codes, building on a previous proposal of hyperinvariant tensor networks. Its tensor network description consists of local, non-perfect tensors describing CSS codes interspersed with Hadamard gates, placed on a hyperbolic $\{p,q\}$ geometry with even $q\geq 4$, yielding an infinitely large class of subsystem codes. We construct an example for a $\{5,4\}$ manifold and describe strategies of logical gauge fixing that lead to different rates $k/n$ and distances $d$, which we calculate analytically, finding distances which range from $d=2$ to $d\sim n^{2/3}$. Investigating threshold performance under erasure, depolarizing, and pure Pauli noise channels, we find that the code exhibits a depolarizing noise threshold of about $19.1\%$ in the code-capacity model and $50\%$ for pure Pauli and erasure channels under suitable gauges. We also test a constant-rate version with $k/n=0.125$, finding excellent error resilience (about $40\%$) under the erasure channel. Recovery rates for these and other settings are studied both under an optimal decoder as well as a more efficient but non-optimal greedy decoder. We also consider generalizations beyond the CSS tensor construction, compute error rates and thresholds for other hyperbolic geometries, and discuss the relationship to holographic bulk/boundary dualities. Our work indicates that Evenbly codes may show promise for practical quantum computing applications.
1 Introduction
@@
In this work, we explore a recent proposal for hyperinvariant tensor network (HTN) codes [24] and show that it encompasses simple qubit-level codes; we dub these new codes Evenbly codes after the author of the original HTN paper [25]. An essential property of these tensor network codes is that they do not require perfect tensors, and can thus be built from gates that mediate less-than-maximal entanglement. Our specific construction incorporates seed tensors from $\llbracket q,1,2\rrbracket_{2}$ CSS codes placed on the vertices of a hyperbolic tiling, along with Hadamard gates on its edges; this can be generalized to any even-$q$ hyperbolic tiling with Schläfli symbols $\{p,q\}$, giving rise to a new, infinite family of holographic subsystem codes. We methodically show that the resulting code structure upholds the isometric constraints typified by the HTN ansatz [26, 25], and additionally show how to partition the logical subspace as a subsystem code. The distance and rate of variants of the code is analyzed, with some achieving rates up to $R=\sqrt{3}/2$ and others exhibiting a distance scaling up to $d(n)\sim n^{2/3}$. Moreover, we investigate the erasure threshold of the zero-rate $\{5,4\}$ Evenbly code using two types of decoders: a quadratically-scaling greedy reconstruction algorithm and an optimal recovery algorithm based on Gaussian elimination, which exhibits a cubic time complexity [9, 27]. We find that the greedy reconstruction algorithm faithfully provides a lower bound for the threshold (approximately $40\%$), whereas the Gaussian-elimination algorithm obtains a threshold of $49.95\pm 0.01\%$ in the $Z$ gauge. A constant-rate $k/n=0.125$ variant of the Evenbly code is tested as well, and it is shown to achieve a threshold of about $44\%$ in the $Y$ gauge for the central logical qubit, and more than $39\%$ in the $Z$ gauge. In both the zero- and constant-rate case, the $X$ gauge corresponds to a product state embedding of the ungauged logical qubits on the boundary, preserving the transversal weight-$2$ gates of the $\llbracket 4,1,2\rrbracket_{2}$ seed code. We also study the effects of depolarizing and pure Pauli noise on the zero-rate $\{5,4\}$ Evenbly code using the integer-optimization decoder from [28], proving the existence of a threshold at $19.1\pm 0.94\%$ for the former, and around $50\%$ for the latter, ostensibly surpassing the well-known zero-rate hashing bound from random coding theory [29, 30, 31]. Other constructions are possible, and we provide several examples that utilize seed tensors which are neither CSS nor planar 2-uniform, as well as evaluating the erasure thresholds under greedy decoding for several even-$q$ generalizations of the Evenbly code discussed in the main text.
@@
This paper is organized as follows. We start with a synopsis on tensor networks and holographic quantum codes (Section 2.1) and subsequently introduce hyperinvariant tensor networks (HTN) as an ansatz (Section 2.2). Next, we construct Evenbly codes (Section 3), commencing first with the introduction of 1-uniform (i.e. GHZ) states to the vertex tensors of the original HTN ansatz (Section 3.1), and following subsequently with the construction of the qubit-level Evenbly code using planar 2-uniform states (Section 3.2). We further show that our construction admits a subsystem structure, in line with previous holographic quantum codes, and we describe how to partition the logical space and perform gauge fixing for the code (Section 3.3). Proceeding, we investigate the rate and distance scaling for the code, and as a result construct a finite-rate version of the Evenbly code (Section 3.4); the threshold properties are investigated for a zero-rate Evenbly code (Section 3.5) and a constant-rate variant in Section 3.6; in the process, we explain how the $X$ gauge choice in the Evenbly code allows for weight-2 transversal logical operations to emerge, and we show that very high thresholds emerge in several of the other gauge choices. Finally we report depolarizing and pure Pauli noise threshold results (Section 3.7), which are found to be competitive with known code-capacity thresholds for topological codes [6]. The implications of our findings are discussed and concluding comments are provided in Section 4. In the appendices, we provide additional information on how to build multipartite maximally-entangled (MME) states (Appendix A); an example construction of an Evenbly code using an AME graph state (Appendix B); Several additional constructions on generalized $\{p,q\}$ tilings and and an accompanying greedy threshold analysis (Appendix C); several seed-tensor constructions that are neither CSS nor planar 2-uniform states, yet still can be used to create new Evenbly codes (Appendix D); and lastly, details on the three different decoder strategies employed in this paper (Appendix E).
2 Background
2.1 Tensor Networks & Holographic Codes
@@
Holographic codes typically exhibit uberholography, i.e., only operators on a fractal subset of the boundary are needed in order to reconstruct logical operators in the bulk [40]. Uberholography is a special case of operator-algebra quantum-error correction [41], in which a subalgebra representing all possible logical operators in a region is defined using non-local physical boundary operators. HaPPY-style holographic codes have been shown to exhibit high thresholds under both erasure and depolarizing noise models [28, 14, 13], as well as exhibiting biased-noise threshold behavior better than the hashing bound [42, 43]. In the spirit of investigating holographic dualities, it has been shown that HaPPY-style holographic codes do not reproduce many of holography’s known features at finite $N$ (i.e., under gravitational corrections) [44, 45]. Such features include non-trivial entanglement spectra of reduced density matrices, state-dependent bulk-reconstruction, and corrections to the entanglement entropy, whose recovery requires tensor network codes with non-perfect tensors [11, 12, 24]. The quantum code construction that we propose in this paper, in addition to its novel error correction properties, are thus relevant from the perspective of holography as well; we address these attributes in our companion paper [24].
2.2 Hyperinvariant Tensor Networks
Figure 1: Codes from a hyperinvariant tensor network (HTN). (a) Evenbly’s HTN setup from [25], composed of $A$ and $B$ tensors on a hyperbolic $\{p,q\}$ tiling (here for $p=5,q=4$) that respect the symmetries of the infinitely extended tiling. Open black edges represent a state on the physical boundary qubits, while connected black edges are tensor network contractions. (b) The extension of the HTN into a quantum error correction code proposed in [24] (dubbed Evenbly code), in which the $q$-leg AA tensors are replaced by q+1q{+}1-leg $A^{\prime}$ tensors, with the additional leg (red dot) representing a logical “bulk” qubit. The tensor network forms the encoding isometry of this bulk-to-boundary code.
@@ 2.2 Hyperinvariant Tensor Networks
Figure 2: Isometry constraints for the HTN defined on a $\{p,q\}$ tiling. We color-code the relevant tensors as follows: light-/dark-gray tensors represent AA and $A^{\dagger}$; light-/dark-blue tensors represent $B$ and $B^{\dagger}$; light-/dark-purple tensors depict the $3:1$ isometries $W$ and $W^{\dagger}$; and the $4:2$ isometries $U$ and $U^{\dagger}$ are shown in creme and light orange, respectively.
3 Results
3.1 HTN Ansätze from 1-Uniform States
We begin by constructing first a simpler case of the original HTN ansatz of [25]; namely, we prepare a single boundary state rather than a bulk-to-boundary code, which permits us to analyze the tensor network structure more easily. We then show in Section 3.2 that this simplified case easily generalizes to a new subclass of holographic quantum codes.
@@ 3.1 HTN Ansätze from 1-Uniform States
    AA and $B$ fulfill isometry constraints involving tensors on single and neighboring vertices. The constraints are shown explicitly in Figure 2 for the $q>3$ case (the special $q=3$ case is covered in [25]).
@@
    AA is symmetric under cyclic index permutations, i.e., $A_{k_{1},k_{2},\dots,k_{q}}=A_{k_{q},k_{1},\dots,k_{q-1}}$.
@@
    AA is chosen as a non-perfect tensor, i.e., there exist bipartitions of its $q$ indices between which it does not mediate maximal entanglement.
@@
[…]
@@
$\mathcal{G}_{1}$ $=X^{\otimes q},$ (2a)
$\mathcal{G}_{k}$ $=I^{\otimes k-2}(ZZ)I^{\otimes q-k}\ ,$ (2b)
where $k\in\{2,3,\dots,q\}$. As usual, we end up with (nk)=q(n-k)=q generators for the full stabilizer group.
@@
$H=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix}\$. (3)
The isometry constraints from the remaining first criterion can be proven by using operator pushing: By applying the stabilizers from Equation 2a, we can replace a Pauli operator acting on one of the tensor legs by an equivalent operator acting on other legs. Moving Pauli operators past a $B$ tensor involves exchanging $X\leftrightarrow Z$, as $XH=HZ$ and $ZH=HX$.
@@
Figure 3: Operator recovery on subsystems of the HTN with $A$ tensors (light gray) chosen to represent GHZ states and $B$ tensors (light blue) chosen as Hadamard gates. A) For the isometry $W$, applying the generators from Equation 2a removes a Pauli operator on the top leg while adding new ones on the bottom ones. As this is possible for the whole Pauli algebra (generated by X and Z), the tensors form an isometry from top to bottom legs, as expected from the 1-uniformity requirement. Note that pushing through a $B$ tensor swaps $X\leftrightarrow Z$. B) For the $U$ isometry, a Pauli operator can again be pushed to the bottom legs by applying the relation from Equation 2a. Due to reflection symmetry, it is sufficient to consider operators acting on only one of the two top legs. These diagrams offer proof that the generalized isometry constraints shown in Figure 2 are upheld, thus allowing for the construction of a consummate HTN.
Figure 4: Operator recovery on subsystems of the Evenbly code (Equation 4) via the isometry constraints. A)-B) For a single $A^{\prime}$ tensor, the logical $\bar{X}$ and $\bar{Z}$ operators (red) can be represented on $\frac{q}{2}$ and 22 sites, respectively, following Equation 6. Note that X and Z operators are swapped when pushed through a $B$ tensor. For two adjacent $A^{\prime}$ tensors and an intermediate $B$ tensor, the $\bar{Z}$ operator on a given logical qubit can be represented on the physical qubits on one side. The $\bar{X}$ operator requires a representation on both sides, with an application of the stabilizers (Equation 4) removing any Pauli operators acting on contracted legs. By reflection symmetry, the same rules apply for logical operators acting on the second site. C)-D) Any operator acting on a single physical leg can also be pushed to the remaining ones by applying Equation 4. Together with A)-B), these rules are sufficient to prove the isometry conditions.
3.2 HTN qubit codes
@@
Now we generalize the results gained from Section 3.1 and define Evenbly codes; i.e., HTNs with bulk logical degrees of freedom. Evenbly codes were first defined in [24] as tensor network codes that associate a logical qubit with each vertex in an HTN, a process of bulk implantation which replaces each $q$-leg AA tensor by a $(q+1)$-leg tensor $A^{\prime}$. In analogy with Definition 1, an Evenbly code is then defined as follows:
@@
Note that $A^{\prime}$ merely describes an error-detecting code with code distance $d=2$; this condition implicitly excludes perfect tensors for $q>4$, which would lead to codes with $d>2$. Though it is possible to define a five-qubit perfect $A^{\prime}$ at $q=4$, the tensor describing any physical $4$-qubit state after projecting out the logical qubit will generally be non-perfect, thus again leading to tensor network states with non-flat entanglement spectra, generalizing the HTN definition of Section 3.1. Generally, finding solutions to the constraints in Figure 5 is much harder than for those in Figure 2 since the isometries $W^{\prime}$ and $U^{\prime}$ act on more input sites than $W$ and $U$.
Ref. [24] introduced only a single ququart example fulfilling these constraints; we will now show that our GHZ HTN construction can be extended to define a class of qubit codes whose $A^{\prime}$ tensors define Calderbank-Steane-Shor (CSS) codes. While we focus on this construction in the remainder of the paper, other qubit-level Evenbly codes can be constructed: In Appendices B and D, we introduce Evenbly codes whose seed tensors are not built from CSS codes, with $A^{\prime}$ tensors describing strictly 2-uniform or AME(5,2)(5,2) states. In contrast, the Evenbly code can be defined for any even $q\geq 4$, and its stabilizer generators defining the $A^{\prime}$ tensors of rank $(q+1)$ are given by
@@
[…]
The tensor $A^{\prime}$ is therefore partially 2-uniform when considering the logical and any one physical index, fulfilling the first of the Evenbly code constraints. In fact, it fulfills an even stronger condition: For any planar embedding of the logical index, the tensor describes a q+1q{+}1-qubit state that is planar 2-uniform, i.e., any two qubits that are neighboring under this embedding are maximally entangled with the remainder [52]. One choice for the logical operators is given by
@@
3.3 Evenbly Codes as Subsystem Codes